# Integration of TinyML and LargeML: A Survey of 6G and Beyond

### v1

@@ II-B2 LargeML Operation
LargeML operates through a multi-stage process, beginning with comprehensive data collection and preprocessing to ensure quality. It undergoes initial training on advanced algorithms and high-performance hardware to identify data patterns, followed by fine-tuning for specific tasks to enhance accuracy. Post-deployment, LargeML processes real-time user inputs, delivering precise responses while continuously adapting to feedback and new data, maintaining its accuracy and relevance over time.
However, LargeML operation varies with the application domain and character of input data from sensors and devices [5]. In wireless communication, LargeML is typically used to optimize connectivity and manage network traffic across mobile networks, Wi-Fi systems, and satellite links. In IoT contexts, LargeML analyzes data from smart home devices, connected vehicles, and healthcare monitors to support intelligent automation and real-time decisions. In smart homes, LargeML functions as an AI control platform, adjusting lighting, temperature, and security systems based on user behaviors. In connected vehicles, LargeML enables real-time navigation, vehicle diagnostics, and autonomous driving by processing sensors and wireless network data. In healthcare, LargeML supports remote monitoring of vital signs, enabling prompt interventions and personalized care.
LargeML can also be deployed on edge devices and cloud platforms for efficient system management [52]. Edge deployment reduces latency by enabling local data processing, while cloud integration offers the scalability required to manage large volumes of data and user interactions. Continuous monitoring and updates based on user feedback and new data help maintain the model’s accuracy and relevance. Such integration across wireless and IoT systems enables more responsive and efficient solutions in a range of domains. For example, LargeML can optimize traffic flow and energy consumption in smart city infrastructure, enable predictive maintenance through industrial IoT monitoring, and improve agricultural practices based on real-time environmental data.
@@ II-B3 Performance Evaluation
Given its broad application scope, the performance of LargeML can be evaluated at three levels. At the learning level, key indicators include model accuracy, output quality, and operational efficiency. Model accuracy is measured using metrics [55] such as precision, recall, F1 score, and area under the receiver operating characteristic (AUC) curve, which assesses the model’s ability to correctly classify inputs. For example, AUC measures the model’s ability to discriminate between classes. Model output quality is assessed using metrics such as perplexity, bilingual evaluation understudy (BLEU) score, and recall-oriented understudy for gisting evaluation (ROUGE). Perplexity gauges the model’s confidence in predicting subsequent outputs. BLEU and ROUGE compare the model’s generated output to reference resources.
At the operational level, performance is measured by efficiency metrics, such as latency, throughput, and resource use. Latency refers to the time required for the model to generate a response. Throughput indicates the number of requests the model can process per unit of time. Resource use tracks the computational resources consumed to ensure the model operates efficiently without overloading the system.
At the service level, evaluation focuses on user experience, ethics, and accountability [17, 19]. User experience is assessed by: (i) User satisfaction to capture user perceptions of model performance via direct feedback, (ii) Engagement to measures the frequency and depth of user interactions with the model, and (iii) Retention rate to indicate the proportion of users who continue using the model over time. Meanwhile, ethics and accountability metrics assess whether LargeML meets critical standards: (i) Fairness to evaluates whether the model’s predictions are fair across different user groups, helping identify and reduce bias, (ii) Transparency to measure how clearly users and stakeholders can understand the model’s decisions, often supported by explainability tools, and (iii) Ethical compliance to ensures adherence to privacy, security, and accountability standards, fostering trust in the system.
@@ III-A Challenges in TinyML
III-A1 Resource Constraints
TinyML devices, such as edge IoT nodes equipped with sensors or microcontrollers, operate under stringent processing power, memory, and energy constraints that complicate the deployment of complex ML models, forcing trade-offs between model accuracy and device capabilities. The resource constraints for TinyML device vary. For example, as noted in [56], microcontrollers may lack DRAM and an operating system and operate with up to 1 MB of read-only flash memory and up to 256 KB of SRAM. The study in [11] compares hardware platforms that support TinyML devices, reporting flash memory capacities of up to to 128 MB and SRAM capacities up to 64 MB. Consequently, given the absence of sufficient CPU, GPU, or TPU resources, the deployment and training of ML models on such devices is impractical. Larger models, although more accurate, often exceed the memory capacities of TinyML devices, and require aggressive model compression or simplification. In the context of 6G, such limitations become more pronounced as the demand for real-time, data-intensive applications increases.
III-A2 Model Compression
The deployment of ML models on TinyML devices requires compression techniques such as pruning, quantization, and knowledge distillation to reduce model size and complexity [57]. However, these methods often degrade model accuracy and performance, introducing the need for a careful trade-off between compression efficiency and model integrity. For example, quantized graph optimization presents particular challenges due to mixed-precision tensors and the absence of batch normalization layers [58]. In addition, TinyML devices lack the computational capacity to support full backpropagation. Although pruning reduces model size by eliminating less critical parameters, it may also inadvertently remove essential parameters, resulting in lower prediction accuracy.
III-A3 Energy Efficiency
Many TinyML devices operate on battery power, requiring models with high energy efficiency to ensure sustained operation. The design of TinyML models that minimize power consumption while satisfying performance requirements is a challenging task, given the computational overhead of ML inference. Poor energy efficiency can quickly deplete power sources, shorten device lifespan, and limit deployment in remote or inaccessible locations. For example, continuous data processing on a battery-powered IoT sensor can rapidly exhaust the power supply if the model lacks an energy optimization mechanism [40]. In real-time applications, such as autonomous vehicles or industrial automation, which are key areas in the 6G paradigm, low-latency inference is critical. Achieving fast inference without sacrificing energy efficiency requires advanced optimization strategies.
III-A4 Data Preprocessing and Limited Memory Resources
TinyML devices face significant constraints in performing on-device data preprocessing due to limited computational and memory resources. Tasks such as normalization, augmentation, and feature extraction require substantial processing power and memory, which TinyML devices generally lack. For example, image data captured by a camera may need extensive preprocessing before it can be used for object recognition [59]. Reduction of the dataset to satisfy memory constraints can also impair model robustness and reliability.
III-A5 Mission-Critical Reliability
TinyML’s limited numerical precision and potential for error during circuit fabrication and wafer-level assembly raise concerns about system reliability. These concerns are amplified in mission-critical applications. For example, in healthcare, TinyML may be used for remote surgery, vital sign monitoring, diagnosis, or treatment management, where any error could lead to severe or life-threatening consequences [7]. Reliability and robustness are therefore non-negotiable in such contexts, where device failures could directly affect patient safety.
III-A6 Data Privacy and Security
TinyML frequently processes data from sensors such as cameras and microphones, which may inadvertently capture sensitive personal information. This raises serious privacy concerns, especially as 6G is expected to interconnect billions of IoT devices[11]. Implementation of privacy-preserving techniques, such as differential privacy or secure multiparty computation, is challenging on TinyML platforms due to their limited computational capacity. Such limitations also increase vulnerability to unauthorized manipulation or access by malicious actors [60].
@@ III-B Challenges in LargeML
III-B1 High Resource Consumption
LargeML models, which contain millions to trillions of parameters [61], require substantial computational resources for both training and inference. These models typically depend on high performance GPUs or specialized hardware such as TPUs. Training LargeML can span days or weeks, consuming considerable energy and raising environmental concerns. Their size, along with huge datasets, necessitates extensive memory and storage capacity. High-capacity RAM and storage systems are essential equipment for managing weights and activation states during both training and inference.
III-B2 Latency Sensitivity and Real-Time Processing
The complexity and size of LargeML models result in increased inference latency. In addition to execution time, these models are often deployed on cloud servers to meet their resource demands, requiring frequent data transfers between edge devices and the cloud. This dependence introduces network latency, which can be inconsistent and difficult to control. Many 6G applications demand real-time or near-real-time responsiveness, requirements that LargeML may struggle to meet. For example, 6G real-time applications such as augmented reality, virtual reality, and autonomous driving are highly sensitive to latency and cannot accommodate the delay caused by both inference processing and data transmission. [62].
III-B3 Deployment and Accessibility
The deployment of LargeML involves a series of complex tasks, including hardware setup, software configuration, and model optimization for specific applications. These steps require substantial technical expertise and financial investment. Hardware, such as GPUs, TPUs, and high-capacity servers, is often prohibitively expensive. For advanced LargeML models designed to achieve high accuracy on complex tasks, the costs can further restrict accessibility, reproducibility, and broader adoption, particularly among small and medium-sized enterprises and independent researchers [63]. In many developing regions, infrastructure to support LargeML remains inadequate, hindered by unreliable internet access, limited data centers, and a shortage of skilled professionals. These disparities widen the digital divide, limiting the reach of AI advances as the 6G era approaches.
III-B4 Scalability
The scaling of LargeML presents significant challenges due to high computational costs, data management complexities, and stringent network requirements [64]. Resource demands frequently exceed the capabilities of current infrastructure. LargeML requires extensive data for training and continuous updates, with consequent resource-intensive, complex data management, compounded by heterogeneous datasets sourced from a range of domains served by 6G networks. Managing such datasets is resource-intensive, and inefficient handling leads to data bottlenecks and degraded model performance. Integration of heterogeneous data sources adds further complexity, requiring extensive preprocessing and validation. As model scale increases, supporting infrastructure must enable automation and orchestration to manage higher data throughput and maintain low-latency communication between distributed systems. Network constraints can become critical bottlenecks, undermining scalability and compromising the real-time performance demanded by 6G applications.
III-B5 Interpretability and Fine-Tuning
LargeML models often function as “closed boxes”, with internal decision-making processes that lack transparency. The complexity of interactions between millions of parameters and layers obscures how specific decisions or predictions are generated. This opacity is unacceptable in mission-critical 6G applications such as healthcare, autonomous vehicles, and finance, where trust and accountability require explainable outputs. Regulatory requirements, for example the European Union’s General Data Protection Regulation [65], explicitly mandate the right to explanation, further reinforcing the need for interpretable AI. Effective model adaptation also depends on fine-tuning that is representative of the target data domain [66]. Fine-tuning involves adjustment of learning rates and batch sizes, regularization of parameters, or retraining of model components. This process must balance performance gains against the risks of overfitting or underfitting problems. Inconsistencies between training and data fine-tuning degrade model accuracy and limit real-world applicability.
III-B6 Data Privacy and Security
LargeML models typically aggregate data from multiple sources into centralized databases to support server-based training and inference. However, such centralization introduces significant privacy risks, as large repositories of sensitive information become prime targets for cyberattack [67]. Even during inference, LargeML models may requires access to confidential data, complicating privacy protection in real time, particularly when deployed across distributed networks. In addition to data privacy, securing LargeML against threats such as model extraction [68] and adversarial attacks [69] remains critical. Attacks can compromise model integrity and expose sensitive training data. For example, an adversary may reverse-engineer a model to recover private information embedded in its training data.
III-C The Vision for 6G: Objective Requirements and the Assistance of Bidirectional Integration
Figure 3: Envisaged 6G capabilities and corresponding pivotal general requirements.
IMT-2030 (6G) will support 15 key capabilities: nine enhanced from 5G (e.g., reliability, latency, data rates) and six new ones (e.g., coverage, sensing, AI), as per [70]. Some metrics are qualitatively described due to complexity [71]. These align with six 6G requirements: ubiquitous connectivity, extreme performance, high intelligence, strong security, green communication, and efficient ISAC (see Fig. 3).
Standalone technologies cannot meet 6G demands (TinyML and LargeML limitations are evident in Sections III-A, III-B), and thus, the focus will shift to an open wireless architecture, with integrations of multiple key technologies from 4G and beyond-5G [72]. Bidirectional TinyML–LargeML integration leverages 6G infrastructure: TinyML processes data at the edge, sending critical information to LargeML for advanced analytics and decision-making. Specific approaches (TL, FTL, SL, FSL) are detailed in Section IV.
@@ III-C1 Ubiquitous Connectivity
6G will accommodate the exponential growth of interconnected devices, including IoT, sensors, wearables, and smart infrastructure, via scalable architecture and dynamic resource management. It must support connection densities of $10^{6}-10^{8}$ devices/km2 and data volumes five times greater than 5G [73]. Key enablers include network slicing and virtualization [74], which allow efficient resource allocation across diverse devices. Deep integration with AI and ML frameworks will enable pervasive, intelligent functionality. A promising approach involves the bidirectional integration of TinyML and LargeML. TinyML handles local data processing at the edge, reducing latency and offloading servers, while LargeML, deployed across a ubiquitous server set (cloud, edge, mobile devices) [75], provides a global view for system-wide optimization. This approach supports decentralized computing and further enhances scalability.
Beyond scalability, 6G aims to deliver ubiquitous global coverage, including terrestrial, aerial, satellite, and underwater systems [76, 77], ensuring connectivity in remote, challenging environments. It must also support high-mobility platforms ($500-1000$ km/h), requiring seamless handovers and latency management strategies [78]. AI-powered techniques like mobility prediction and intelligent handover strategies are expected to ensure seamless and reliable connectivity [79]. The integration of TinyML and LargeML is central to this approach. TinyML enables real-time monitoring and handover prediction at the device level, while LargeML aggregates data across the network to refine mobility and handover strategies. Despite the heterogeneity and non-independent and identically distributed (non-IID) nature of data generated by edge devices [80], LargeML frameworks can facilitate collaborative learning and efficient data aggregation. For example, wearable devices equipped with TinyML can monitor patient vitals and detect anomalies locally. Given the non-IID nature of health data, from among patients and over time from a single patient, LargeML performs server-side aggregation and analysis to support personalized healthcare delivery.
@@ III-C2 Extreme Performance Networking
6G networks require extreme performance across multiple dimensions, including ultra-high peak and user-experienced data rates, ultra-low latency, and high reliability. Quantitatively, reliability targets range from $99.999$ to $99.99999$ with latency goals between $0.1$ and $1$ ms. As outlined in [81], other metrics are qualitative. Data rate expectations for 6G far exceed those of 5G. Peak data rates are projected between $50$ to $200$ Gbps, potentially reaching $1$ Tbps, while user-experienced rates are expected to range from $300$ to $500$ Mbps.
Integrating TinyML and LargeML enhances 6G systems by combining real-time, lightweight local processing with advanced server-side computation. TinyML reduces latency and bandwidth use for simple tasks, while LargeML handles complex computations. In the 6G environment, where servers are ubiquitous, exchanging trained models instead of raw data lowers communication latency, and hierarchical model distribution improves response times and throughput. Bidirectional integration enables collaborative learning through model splitting and transfer, optimizing distributed computing for both edge and server training. From a reliability perspective, integrating TinyML and LargeML harnesses the powerful capabilities of LargeML for server-based decision-making in mission-critical applications. These models can be fine-tuned in real time using data from TinyML devices, ensuring consistent performance and reliability. Integration allows multiple devices to jointly validate shared data without compromising privacy or security. Redundancy in this manner strengthens model accuracy and resilience to data loss or corruption.
@@ III-C3 High Intelligence
The 5G era has evolved alongside rapid developments in AI, propelling progress toward autonomous machine intelligence. With the advent of 6G, a new phase of distributed intelligence is emerging, marked by seamless integration of distributed data processing, learning, computing, and AI inference across large-scale, interconnected systems [81]. This network of smart devices is strengthened by the complementary roles of TinyML and LargeML. TinyML enables edge devices to perform local data processing and training, transmitting only essential outputs to servers, thereby enabling real-time decisions. LargeML, operating on powerful servers, handles resource-intensive tasks, optimizing service intelligence and personalization through a decentralized architecture. Besides, the bandwidth demands of 6G use cases are substantial, for instance, high-accuracy ISAC requires 0.75 GHz, while holographic communications may need up to 1.1 GHz [73]. To meet these needs, AI-driven spectrum optimization will be vital. Real-time AI monitoring can guide dynamic spectrum allocation, reducing congestion and enhancing frequency utilization. AI-empowered cognitive radio networks will enable devices to sense and adapt to their environments, while AI-enabled signal processing will improve signal quality and spectrum efficiency. AI will also facilitate rapid interference mitigation, spectrum sharing, and reallocation of underused bands, ensuring optimal network performance. The convergence of AI and 6G is poised to drive innovation across domains such as smart cities, autonomous transport, healthcare, and immersive entertainment. AI-integrated 6G networks will support real-time urban infrastructure management, enhance the safety and responsiveness of autonomous vehicles, expand access to AI-powered medical diagnostics and remote care, and deliver deeply interactive, high-fidelity entertainment experiences.
III-C4 High Privacy and Security
Privacy and security are critical in deploying ubiquitous AI tasks over 6G networks [82]. As the volume of sensitive data grows, protecting it from unauthorized access and addressing ethical concerns, such as transparency in AI decision-making and privacy risks, becomes imperative. Increasing interconnectivity and system complexity heighten vulnerability to cyber threats, necessitating robust security frameworks that comply with evolving data protection regulations without compromising performance.
While the integration of TinyML and LargeML helps mitigate some privacy concerns, further technologies are needed to address security comprehensively. TinyML enables local preprocessing at the edge, extracting essential features and minimizing the transmission of sensitive data, thus preserving user privacy. It also supports collaborative model training without sharing raw data, reducing breach risks. However, TinyML’s limited resources may restrict implementation of advanced privacy-preserving techniques. Centralized and distributed data repositories remain key targets for attacks in a hyperconnected 6G landscape. TinyML further enhances resilience by enabling real-time anomaly detection at the edge, allowing for rapid threat identification and response. In tandem, LargeML systems leverage high-performance servers to perform in-depth threat analysis using aggregated data from multiple devices. These systems can deploy adaptive, sophisticated security mechanisms, as discussed in [83], bolstering the overall security posture of 6G networks.
@@ III-C5 Green Communications
Green communications prioritize energy efficiency and sustainability, core elements of the 6G vision aimed at reducing CO2 emissions and minimizing the environmental impact of communication technologies [84]. As 6G expands the number of connected devices, energy consumption is expected to rise, making the reduction of the information and communication technology sector’s carbon footprint a critical goal. Energy-efficient systems not only cut operational costs but also benefit both providers and users. The combined use of TinyML and LargeML advances this goal by optimizing resource allocation and energy efficiency across the network. TinyML enables low-power devices to process data locally, reducing transmission frequency and reliance on energy-intensive servers. LargeML complements this by efficiently managing data at the server level, leveraging its computational capacity without unnecessary overhead. Their integration distributes computation across edge and cloud environments, alleviating the load on central servers and data centers, thereby significant reducing CO2 emissions.
TinyML further supports sustainable operations by adapting to real-time energy constraints and incorporating energy harvesting (e.g., solar or kinetic sources), minimizing dependence on traditional power. When paired with LargeML, these systems maintain performance even under tight energy budgets. In environmental monitoring, TinyML enables on-site anomaly detection, forwarding only critical data to centralized systems for deeper analysis, facilitating timely, informed responses. Additionally, their bidirectional integration supports predictive maintenance of network infrastructure, identifying issues early to prevent energy inefficiencies and environmental degradation. Such proactive measures are vital to reducing the ecological footprint of future communication networks.

### v2

@@ II-B2 LargeML Operation
LargeML follows a multi-stage process: large-scale data collection and pre-processing, initial training on high-performance hardware, fine-tuning for specific tasks, and post-deployment adaptation through real-time inputs and feedback. This ensures both accuracy and long-term relevance. Its operation varies across domains and input data characteristics from sensors and devices [193]. In wireless communication, LargeML optimizes connectivity and manages network traffic across mobile, Wi-Fi, and satellite links. In the IoT domain, it processes data from smart homes, connected vehicles, and healthcare monitors to enable intelligent automation and real-time decision-making. For example, it serves as an AI control hub in smart homes, supports navigation and diagnostics in vehicles, and enables remote monitoring and personalized care in healthcare. Deployment strategies include edge-based processing to reduce latency and cloud integration for scalability [60]. Continuous monitoring and updates based on user feedback and new data sustain model performance across applications. Integrated into wireless and IoT systems, LargeML supports smart city infrastructure (e.g., traffic and energy optimization), predictive industrial maintenance, and smart agriculture through environmental monitoring.
@@ II-B3 Performance Evaluation
The performance of LargeML can be assessed at three levels, described below.
At the learning level, the focus is on accuracy, output quality, and efficiency. Accuracy is assessed using precision, recall, F1 score, and area under the receiver operating characteristic (AUC), which measure classification and discrimination ability [4]. Output quality is gaged using perplexity (confidence in predictions), bilingual evaluation understudy (BLEU), and recall-oriented understudy for gisting evaluation (ROUGE), which compare generated text against reference outputs.
At the operational level, performance is measured by typical metrics, such as latency (response time), throughput (requests processed per unit time), and resource use (computational cost), ensuring scalable and reliable operation.
At the service level, evaluation focuses on user experience, ethics, and accountability [24, 207]. User experience includes user satisfaction (perception feedback), engagement (interaction depth), and retention rate (continued usage’s proportion). Ethical and accountability metrics assess fairness (bias reduction), transparency (explainable decisions), and ethical compliance (privacy, security, and accountability standards).
@@ III-A Challenges in TinyML
III-A1 Resource Constraints and Energy Efficiency
TinyML devices typically operate with limited processing power, memory, and energy resources, complicating the deployment of complex models. For instance, microcontrollers often lack DRAM and an operating system, with memory capacities ranging from a few hundred kilobytes to several megabytes [96, 1]. Such limitations preclude the use of LargeML without aggressive compression, often forcing trade-offs between accuracy and feasibility. Furthermore, many TinyML devices are battery-powered, requiring energy-efficient models that minimize inference overhead to extend device lifespan. Achieving low-latency inference while maintaining energy efficiency remains a key challenge, especially for mission-critical 6G applications such as autonomous driving and industrial automation [169].
III-A2 Model Deployment and Data Pre-processing
The deployment of ML models on TinyML devices necessitates compression techniques such as pruning, quantization, and knowledge distillation [154]. Although these approaches reduce memory and computation costs, they may compromise accuracy, particularly when parameters essential for prediction are removed. Quantized optimization introduces additional difficulties due to mixed-precision tensors and the absence of batch normalization layers [68]. Beyond model compression, pre-processing tasks such as normalization, augmentation, and feature extraction require computational and memory resources that are typically unavailable on TinyML devices. This makes handling complex datasets, such as high-resolution images, particularly challenging.
III-A3 Reliability, Privacy, and Security
TinyML devices suffer from limited numerical precision and fabrication‑related errors, which can undermine system reliability, especially in mission‑critical settings like healthcare, where failures may endanger patient safety [167]. They also face significant privacy and security risks: although they often handle sensitive audio or visual data, their constrained compute capacity prevents the use of stronger protections, such as differential privacy or secure multiparty computation, leaving them more exposed to data leakage and adversarial attacks [149].
@@ III-B Challenges in LargeML
III-B1 Resource and Infrastructure Demands
Training and inference of LargeML require specialized hardware such as GPUs or TPUs, along with extensive RAM and storage for managing parameters and activation states [81]. Training often spans days or weeks, incurring high financial and energy costs. Such requirements limit accessibility, particularly for small and medium enterprises or researchers in developing regions, where infrastructure and expertise may be insufficient [52].
III-B2 Latency and Scalability Issues
Due to their size and complexity, LargeML models are prone to high inference latency. Offloading computation to cloud servers introduces further network delays, which are unacceptable for latency-sensitive 6G applications such as augmented reality, virtual reality, and autonomous vehicles [117]. Scaling LargeML is further hindered by the need to manage vast, heterogeneous, and often non-independent and identically distributed (non-IID) datasets across distributed infrastructures. This increases the risks of bottlenecks, degraded model performance, and challenges in real-time processing [13].
III-B3 Deployment, Interpretability, and Adaptability
Deploying LargeML involves costly and complex system integration, including hardware configuration, software optimization, and application-specific tuning. Moreover, these models often operate as “closed boxes,” with opaque decision-making processes. This lack of interpretability undermines trust in mission-critical domains such as healthcare and finance, where explainability and accountability are essential [138]. Effective fine-tuning is also challenging: while adaptation to new data domains is necessary, it requires careful balancing to avoid overfitting or underfitting [39].
III-B4 Privacy and Security Concerns
LargeML models typically rely on centralized data aggregation, which concentrates sensitive information in repositories vulnerable to cyberattacks [38]. Inference tasks may also expose confidential data, complicating real-time privacy protection. Furthermore, these models are susceptible to threats such as adversarial attacks, model extraction, and reverse engineering, which can compromise both model integrity and the confidentiality of training data.
III-C The Vision for 6G
IMT-2030 framework identifies six overarching requirements for 6G networks: ubiquitous connectivity, extreme performance, high intelligence, strong security, green communication, and efficient ISAC (see Fig. 3) [105]. Standalone approaches (whether TinyML or LargeML) are insufficient to satisfy these requirements, necessitating a bidirectional integration paradigm. TinyML supports local data processing at the edge, while LargeML leverages high-capacity infrastructure for global-scale optimization. Together, they enable complementary capabilities aligned with the 6G vision.
Figure 3: Envisaged 6G capabilities and pivotal requirements.
@@ III-C1 Ubiquitous Connectivity
6G must support ultra-dense networks, with connection densities up to $10^{6}-10^{8}$ devices/km2 and data volumes five times higher than 5G [104]. TinyML enables edge devices to process local data in real time, reducing network load, while LargeML, deployed across a ubiquitous server set (cloud, edge, mobile devices) [102], aggregates distributed information to optimize resource allocation and mobility management. This integration supports global coverage across terrestrial, aerial, satellite, and underwater systems, with intelligent handover strategies for high-mobility platforms [33, 34].
@@ III-C2 Extreme Performance Networking
Target metrics for 6G include latencies as low as 0.1 ms and reliability levels up to 99.99999, with peak data rates projected to reach 1 Tbps [70]. Bidirectional integration enhances performance by reducing latency at the edge (TinyML) while maintaining complex server-side analytics (LargeML). Model splitting, model transfer, hierarchical distribution, and lightweight fine-tuning enable real-time collaboration across devices and servers, improving throughput, resilience, and accuracy.
@@ III-C3 High Intelligence
6G envisions distributed intelligence across networks, combining localized processing and global decision-making [70]. TinyML provides real-time inference for edge devices, while LargeML ensures resource-intensive optimization and personalization. This collaboration supports bandwidth-heavy applications such as ISAC and holographic communication [104], where AI-driven spectrum allocation and cognitive technologies improve efficiency and responsiveness.
III-C4 Strong Privacy and Security
As 6G interconnects billions of devices, ensuring data confidentiality and system robustness is paramount. TinyML enhances privacy by performing local pre-processing and transmitting only essential features, thereby reducing exposure of raw data. Concurrently, LargeML systems, supported by high-performance servers, can conduct sophisticated threat detection and adaptive defense using aggregated data from multiple devices [67].
@@ III-C5 Green Communications
With the proliferation of connected devices, energy efficiency is central to sustainable 6G systems. TinyML reduces reliance on energy-intensive servers by enabling low-power, on-device processing. LargeML complements this by optimizing network-wide resource management, minimizing redundant processing, and reducing overall CO2 emissions [64]. Bidirectional integration also supports predictive maintenance of network infrastructure and the use of energy-harvesting technologies.